OpenStudy (anonymous):

Write an expression to describe a rule for the sequence. Then find the 100th term in the sequence. 5, 13, 21, 29, 37, 45, … • 8n; 800 • 8n – 3; 797 • 3n – 8; 292 • n + 8; 108 my answer is D check my work plz

OpenStudy (anonymous):

See, here you are asked to find the general term first and using that you are to find 100th term.. Now, in this formula: a n =a+(n−1)d a n is your general term or you can say the nth term. a is your first term in the sequence, (here it is 4) n is number of terms. d is common difference. (In this case : 5) So, these values you have to substitute in that formula, you will get: a n =4+(n−1)(5)⟹a n =4+5n−5⟹a n =5n−1 So, this is your general term.. Now can you find 100th term here? Hint : you have to just put n = 100 in the following general term you have got: a n =5n−1 Here just put n=100 and find a n . General term means a n that in turn means "Rule for a Particular Sequence". Here : a n =5n−1 is Rule of the sequence you are given with.. Hope this helps, have a nice day. Tag me on anymore math help you need!